two complementary angles are in the ratio 4:11. find the measure of the smaller angle
Found 2 solutions by erica65404, stanbon:
Answer by erica65404(394)

complementary angles add up to 90 degrees.
the ratio is 4:11.
There are a total of 15 ratio points. we want to find out how many degrees each point is worth.
we take 90 degrees and divide by 15 and we get 6.
we multiply the ratio 4:11 by 6.
24:66
24 and 66 are the degrees of the angles.

Answer by stanbon(75887)

two complementary angles are in the ratio 4:11. find the measure of the smaller angle
-----
angle:: x
complement:: 90-x
--------
Equation:
x/(90-x) = 4/11
---------
11x = 4(90-x)
11x = 360-4x
15x = 360
x = 24 degrees
90-x = 66 degrees
=====================
Cheers,
Stan H.